# Jailbreak Large Vision-Language Models Through Multi-Modal Linkage

### v3

@@ Abstract
With the significant advancement of Large Vision-Language Models (VLMs), concerns about their␣potential misuse and abuse have grown rapidly. Previous studies have highlighted VLMs’ vulnerability to jailbreak attacks, where carefully crafted inputs can lead the model to produce content that violates ethical and legal standards. However, existing methods struggle against state-of-the-art VLMs like GPT-4o, due to the over-exposure of harmful content and lack of stealthy malicious guidance. In this work, we propose a novel jailbreak attack framework: Multi-Modal Linkage (MML) Attack. Drawing inspiration from cryptography, MML utilizes an encryption-decryption process across text and image modalities to mitigate over-exposure of malicious information. To align the model’s output with malicious intent covertly, MML employs a technique called “evil alignment”, framing the attack within a video game production scenario. Comprehensive experiments demonstrate MML’s effectiveness. Specifically, MML jailbreaks GPT-4o with attack success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench and 99.07% on HADES-Dataset. Our code is available at https://github.com/wangyu-ovo/MML.
@@ 1 Introduction
The rapid development of large vision-language models (VLMs) (Bai et al., 2023; OpenAI, 2024b; Anthropic, 2024) has brought remarkable advancements. Models like GPT-4o demonstrate impressive capabilities in areas such as visual storytelling, character design, and image-based multi-turn dialogues. However, such advancement also raises significant concerns, as the potential misuse of these models could lead to serious consequences.
@@
Figure 1: Comparison of MML with previous structure-based attacks. (a) Existing Structure-based Methods (Gong et al., 2023; Liu et al., 2024c). These methods over-expose malicious content in the input image, such as harmful typographic prompts or elements, along with neutral text guidance, which renders them ineffective against advanced VLMs. (b) Overview of MML using word replacement as an example. When MML converts malicious queries into typographic images, it encrypts them to prevent overexposure of malicious information. During the model inference phase, MML guides the model to decrypt the input and align the output with the malicious intent.
@@
Jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a) pose a huge security concern for Large Language Models (LLMs) and have become a focus of recent research. Despite having undergone safety alignment training (Ouyang et al., 2022; Bai et al., 2022) prior to deployment, most of the attacks can still exploit carefully designed inputs to bypass these safeguards, prompting the models to generate harmful content misaligned with human values. Extensive research has been conducted on jailbreak vulnerabilities (Zou et al., 2023; Wei et al., 2023a; Li et al., 2024a; Chao et al., 2023) and defenses (Lin et al., 2024; Mo et al., 2024; Wei et al., 2023b; Inan et al., 2023) for LLMs, which has in turn sparked similar investigations into VLMs (Dong et al., 2023; Qi et al., 2024; Shayegani et al., 2024; Niu et al., 2024; Gong et al., 2023; Liu et al., 2024c; Li et al., 2024b; Zhang et al., 2023).
@@
Jailbreak attacks on VLMs can be categorized into three types: perturbation-based, structure-based and their combinations. Perturbation-based attacks (Dong et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024) are based on the concept of adversarial examples (Szegedy, 2013; Goodfellow, 2014), employing gradients to create adversarial images. In contrast, structure-based attacks (Gong et al., 2023; Liu et al., 2024c) bypass VLM defenses by embedding harmful content directly into images using typography or text-to-image tools. Li et al. (2024b) propose a method combining perturbation and structure attacks. Their approach amplifies the toxicity of images containing harmful content by applying gradient-based perturbations. Since the most advanced models are largely closed-source, structure-based attacks hold greater potential. This work focuses on structure-based attacks, which are less explored in the literature.
@@
Figure 2: Illustration of MML’s image inputs. Similar to FigStep (Gong et al., 2023), MML converts the malicious query into a typographic image. To prevent direct exposing of harmful information, MML encrypts the input image.
@@
Although current methods have jailbroken LLaVA (Liu et al., 2024a), MiniGPT-4 (Chen et al., 2023), and CogVLM (Wang et al., 2023) with high success rates, their performance degrades when applied to state-of-the-art VLMs such as GPT-4o. This decline can be attributed to two primary limitations of existing structure-based attack methods: over-exposure and neutral text guidance, as shown in Figure 1(a).
@@
Over-exposure occurs when harmful content, such as images of bombs or malicious text embedded in typography, is exposed directly in the input. With advancements in image comprehension and safety alignment, such overt content is likely to trigger rejection by state-of-the-art VLMs.
@@
Neutral text guidance, on the other hand, refers to the absence of stealthy text prompts instructing the model to produce malicious outputs. Consequently, even when the model does not directly refuse to respond, its outputs are often limited to ethical advice, legal reminders, or warnings against harmful behavior—effectively constituting an implicit rejection. An example of this implicit rejection can be found in Appendix A.1.
@@
To address these issues, this work proposes a novel jailbreak attack framework for VLMs, Multi-Modal Linkage (MML) Attack . To mitigate the issue of directly exposing malicious information (over-exposure), MML leverages the linkage between modalities to incorporate an “encryption-decryption” process into the attack, inspired by cryptography. Specifically, MML first “encrypts” images containing harmful information through techniques such as word replacement or image transformation (Figrue 2). It then guides the target VLMs to “decrypt” the original malicious content via text prompts during inference (Figure 3). To counter the lack of malicious guidance, MML combines a technique named “evil alignment”, which describes a virtual scene to align target VLMs’ outputs with malevolent objectives. Figure 1 provides an overview of MML and how it differs from other methods.
@@
To validate the effectiveness of MML attack, we conduct a series of systematic experiments on three benchmarks: SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024c) and HADES-Dataset Li et al. (2024b), and multiple sota large VLMs as target model. The experimental results demonstrate the superiority of our approach, achieving high attack success rates across these datasets. For instance, we successfully jailbreak GPT-4o with success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench and 99.07% on HADES-Dataset. Compared with the baseline methods (Gong et al., 2023; Liu et al., 2024c; Li et al., 2024b), our attack success rate increased by 64.80% ,73.56% and 95.07% respectively.
In summary, the contributions of this work are as follows:
@@
  • We propose Multi-Modal Linkage (MML) Attack , a novel jailbreak attack framework , which draws on cryptography and incorporates an “encryption-decryption” strategy to effectively jailbreak VLMs.
@@
  • We combine MML with a technique called Evil Alignment to amplify the maliciousness of the target model’s outputs. By crafting a virtual scenario about video game production, we effectively aligns the model’s responses with malicious intent.
@@
  • We validate the effectiveness of our approach through extensive experiments, achieving high attack success rate in jailbreaking state-of-the-art VLMs.
@@ 2 Related Work
Jailbreak attack on LLMs.
Jailbreak attacks seek to circumvent safeguards by crafting specific inputs that prompt the LLM to generate harmful content, such as instructions for making a bomb. Early jailbreak methods were manually designed using expert knowledge (Liu et al., 2023). Wei et al. (2023a) argue that jailbreaks occur due to competing objectives and mismatched generalization. Zou et al. (2023) introduced Greedy Coordinate Gradient (GCG), an optimization-based jailbreak attack algorithm, which leverages gradient information from a loss function at the one-hot input layer to generate adversarial suffixes. Liu et al. (2024b) proposed AutoDAN, employing hierarchical genetic algorithms to generate more covert adversarial prompts. Chao et al. (2023) demonstrated the use of LLMs for jailbreak attacks via multi-turn dialogues.
@@
As research into jailbreak attacks on VLMs progresses, evaluating their robustness against jailbreak attacks has emerged as a significant concern. Zhao et al. (2024) pioneered the adversarial robustness evaluation of VLMs. Gong et al. (2023) introduced a benchmark called Safebench, which comprises 500 malicious typographies organized into 10 categories. Liu et al. (2024c) developed MM-SafetyBench, a benchmark featuring 5,040 text-image pairs across 13 scenarios. Additionally, Li et al. (2024b) compiled the HADES-Dataset containing 750 harmful text-image pairs across 5 scenarios. Furthermore, Luo et al. (2024) proposed a more comprehensive benchmark, JailBreakV-28K, which offers enhanced diversity and quality in harmful queries across 16 scenarios. Since this work focuses specifically on structure-based attack evaluation, we select Safebench, MM-SafetyBench and HADES-Dataset as the datasets for our experiments.
@@ 3 Threat Model
Since most of the state-of-the-art VLMs are only accessible via APIs, we follow the black-box attack framework (Gong et al., 2023), which focuses on black-box attacks. Under our setting, the attacker has no knowledge of the target model’s parameters or architecture and can perform attacks only in a single round of dialogue without prior context. The attacker is limited to querying the model and adjusting a few restricted hyper-parameters, such as the maximum token limit and temperature. Notably, we do not alter or introduce any system messages in all experiments.
@@ 4.2.1 Encryption
To reduce the exposure of malicious content, we employ four strategies to encrypt images: word replacement, image mirroring, image rotation, and base64-encoding.
@@
The base64-encoding method (Wei et al., 2023a; Handa et al., 2024), is to first encode the malicious query in base64 format then transform it into a typographical image.
@@ 4.3 Evil Alignment
Another limitation of existing methods is the lack of stealthy malicious guidance in text prompt. Since text safety alignment is more refined (Gong et al., 2023; Liu et al., 2024c), current structure-based attacks rely solely on neutral text prompts to guide the model in generating content related to images containing malicious information (Figure 1(a)). As a result, even when the model doesn’t directly refuse to respond, it often fails to generate the desired malicious content. Instead, it typically provides legal or ethical warnings and suggestions on how to avoid harmful behavior. Our experiments (Section 5.3) also confirm this.
@@
We refer to this as Evil Alignment and we empirically find it has a strong complementary effect to the encrypt-decrypt process. We include the complete prompts in Appendix A.3.
@@ 5.1 Setup
Table 1: Attack Success Rate (ASR) of baseline methods and MML (ours). FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024c). MML-$X$ represents different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. Best results are highlighted in bold. All evaluations are conducted without any system prompt.
@@
We conduct the experiments on three datasets: SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024c) and HADES-Dataset (Li et al., 2024b), which are widely used as benchmarks for structure-based attacks. SafeBench includes 10 AI prohibited topics, selected based on the OpenAI Usage Policy (OpenAI, 2024d) and Meta’s Llama-2 Usage Policy (Meta, 2023). 50 malicious queries are generated by GPT-4 (OpenAI et al., 2024) for each topic, a total of 500 queries. MM-SafeBench comprises malicious queries across 13 scenarios. We filter out no-violation queries by using GPT-4o as moderation, getting a subset of 1,180 queries. HADES-Dataset (Li et al., 2024b) contains 750 malicious instructions across five scenarios. Further details can be found in Appendix A.4.
@@
We set FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024c) as baseline methods for SafeBench and MM-SafeBench. For HADES-Dataset, we use HADES (Li et al., 2024b) as the baseline method. These methods all represent state-of-the-art structure-based or combination-based attacks.
QueryRelated first uses GPT-4 to generate malicious queries in different scenarios and then rewrites these queries. Next, QueryRelated uses GPT-4 to extract unsafe keywords from queries and generates three types of images: Stable Diffusion (SD) images, typography images, and SD + typography images. According to the results from Liu et al. (2024c), SD + typography images method is the state-of-the-art, which is tested as the baseline of our experiments.
HADES uses diffusion models (Chen et al., 2024) to generate malicious images based on harmful instructions. During this process, it employs LLMs (OpenAI et al., 2024) to amplify the harmfulness of the generated images. When targeting white-box VLMs, HADES leverages gradients to generate adversarial noise, enhancing the attack. Finally, the adversarial noise, the malicious image, and the keyword typography are concatenated into a single image in sequence. Since adversarial attacks are transferable (Dong et al., 2023; Zhao et al., 2024), we use the images generated by HADES on LLava-v1.5-7B (Liu et al., 2024a) as the input for the baseline.
Data preprocessing.
@@
Previous methods have demonstrated high jailbreak success rates on VLMs such as LLaVA-1.5 (Liu et al., 2024a) and MiniGPT-4 (Chen et al., 2023), but they struggle with more advanced models like GPT-4o (OpenAI, 2024b). Therefore, we focused on targeting these state-of-the-art models. The models included in our experiment are: GPT-4o-2024-08-06 (OpenAI, 2024b), GPT-4o-Mini-2024-07-18 (OpenAI, 2024a), Qwen-VL-Max-0809 (Bai et al., 2023), and Claude-3.5-Sonnet-20241022 (Anthropic, 2024), all of which represent the most advanced VLMs currently available (Contributors, 2023). The temperature of all models is set to 0.7, and we do not introduce any system messages for all experiments.
@@
Figure 4: ASR of baselines vs. MML-M (ours) across various topics in SafeBench. The left two figures presents the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024c), while the right figure illustrates the ASR of MML using image mirroring as encryption method.
Figure 5: ASR of baselines vs. MML-M (ours) across various topics in $M$-SafeBench. The left two figures presents the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024c), while the right figure illustrates the ASR of MML using image mirroring as encryption method.

### v4

@@ Abstract
With the significant advancement of Large Vision-Language Models (VLMs), concerns about their␣potential misuse and abuse have grown rapidly. Previous studies have highlighted VLMs’ vulnerability to jailbreak attacks, where carefully crafted inputs can lead the model to produce content that violates ethical and legal standards. However, existing jailbreak methods struggle against state-of-the-art VLMs like GPT-4o, due to the over-exposure of harmful content and lack of stealthy malicious guidance. In this work, we propose a novel jailbreak attack framework: Multi-Modal Linkage (MML) Attack. Drawing inspiration from cryptography, MML utilizes an encryption-decryption process across text and image modalities to mitigate over-exposure of malicious information. To align the model’s output with malicious intent covertly, MML employs a technique called “evil alignment”, framing the attack within a video game production scenario. Comprehensive experiments demonstrate MML’s effectiveness. Specifically, MML jailbreaks GPT-4o with attack success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench and 99.07% on HADES-Dataset. Our code is available at https://github.com/wangyu-ovo/MML.
@@ 1 Introduction
The rapid development of large vision-language models (VLMs) (Bai et al., 2023; OpenAI, 2024b; Anthropic, 2024) has brought remarkable advancements. Models like GPT-4o demonstrate impressive capabilities in areas such as image understanding (Zhang et al., 2024) and autonomous driving (Tian et al., 2024). However, such advancement also raises significant concerns, as the potential misuse of these models could lead to serious consequences (Dong et al., 2023; Gong et al., 2023).
@@
Figure 1: Comparison of MML with previous structure-based attacks. (a) Existing structure-based attacks (Gong et al., 2023; Liu et al., 2024b) over-expose malicious content in the input images, such as harmful typographic prompts or elements, along with neutral text guidance, which renders them ineffective against advanced VLMs. (b) Overview of MML attacks. MML first converts malicious queries into typographic images (using word replacement as an example in the illustration) to prevent overexposure of malicious information. In the inference phase, MML guides the model to decrypt the input and align the output with the malicious intent.
@@
Jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a) pose a huge security concern for Large Language Models (LLMs) and have become a focus of recent research. Despite having undergone safety alignment training (Ouyang et al., 2022; Bai et al., 2022) prior to deployment, most of the attacks can still exploit carefully designed inputs to bypass these safeguards, prompting the models to generate harmful content misaligned with human values. Extensive research has been conducted on jailbreak vulnerabilities (Zou et al., 2023; Wei et al., 2023a; Chao et al., 2023; Li et al., 2024a) and defenses (Wei et al., 2023b; Inan et al., 2023; Lin et al., 2024; Mo et al., 2024) for LLMs, which has in turn sparked similar investigations into VLMs (Dong et al., 2023; Gong et al., 2023; Zhang et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024; Liu et al., 2024b; Li et al., 2024b).
@@
Jailbreak attacks on VLMs can be categorized into three categories: perturbation-based, structure-based, and their combinations. Perturbation-based attacks (Dong et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024) are based on the concept of adversarial examples (Szegedy, 2013; Goodfellow, 2014), employing gradients to create adversarial images. In contrast, structure-based attacks (Gong et al., 2023; Liu et al., 2024b) bypass VLM defenses by embedding harmful content directly into images using typography or text-to-image tools. Li et al. (2024b) propose a method combining perturbation and structure attacks. Their approach amplifies the toxicity of images containing harmful content by applying gradient-based perturbations. Since the most advanced models are largely closed-source, structure-based attacks hold greater potential. This work focuses on structure-based attacks, which are less explored in the literature.
@@
Figure 2: Illustration of MML’s image inputs. MML follows FigStep (Gong et al., 2023) to converts the malicious query into a typographic image. But differently, MML encrypts the input image via different methods to prevent direct exposure of harmful information.
@@
Although current methods have jailbroken LLaVA (Liu et al., 2024a), MiniGPT-4 (Chen et al., 2023), and CogVLM (Wang et al., 2023) with high success rates, their performance degrades when applied to state-of-the-art VLMs such as GPT-4o. We suggest the decline might be attibuted decline can be attributed to two primary limitations of existing structure-based attack methods: over-exposure of harmful content and neutral text guidance, as shown in Figure 1(a).
@@
Over-exposure of harmful content occurs when harmful content, e.g. images of bombs or malicious text embedded in typography, is exposed directly in the input. With advancements in image comprehension capability and safety alignment of VLMs, such overt content is likely to trigger rejection.
@@
Neutral text guidance refers to the absence of stealthy text prompts that instruct models to produce malicious and informative outputs while bypassing refusal. Consequently, even when the model does not directly refuse to respond, its outputs are often limited to ethical advice, legal reminders, or warnings against harmful behavior—effectively constituting an implicit rejection. Examples of the implicit rejection are in Appendix A.1.
@@
To address these issues, we propose a novel jailbreak attack framework for VLMs, Multi-Modal Linkage (MML) Attack. MML applies an “encryption-decryption” scheme to the linkage between modalities, which we view as a weak spot of VLMs, to mitigate the over-exposure issue. Specifically, MML first “encrypts” images containing harmful information through different methods including word replacement or image transformation (Figure 2). It then guides the target VLMs to “decrypt” the original malicious content via text prompts during inference (Figure 3). To counter the lack of malicious guidance, MML combines a technique named evil alignment (Zeng et al., 2024), which describes a virtual scene to align target VLMs’ outputs with malevolent objectives. Figure 1 provides an overview of MML and how it differs from other methods.
@@
To validate the effectiveness of an MML attack, we conduct experiments on four latest large VLMs as target models on three benchmarks, i.e., SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b), and HADES-Dataset Li et al. (2024b). The results demonstrate the superiority of our approach, achieving high attack success rates across datasets. For instance, we successfully jailbreak GPT-4o with success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset. Compared with the state-of-the-art baseline methods (Gong et al., 2023; Liu et al., 2024b; Li et al., 2024b), our attack success rate increased by 64.80%, 73.56%, and 95.07% respectively.
In summary, our contributions are as follows:
@@
  • We propose the Multi-Modal Linkage (MML) attack, a novel jailbreak attack framework that draws on cryptography incorporating an “encryption-decryption” strategy.
@@
  • We combine MML with evil alignment to amplify the maliciousness of the target model’s outputs. By crafting a virtual scenario, we effectively align responses with malicious intent.
@@
  • We validate the effectiveness of our approach through extensive experiments on four models and three datasets, achieving a high attack success rate in jailbreaking state-of-the-art VLMs.
@@ 2 Related Work
As research into jailbreak attacks on VLMs progresses, evaluating their robustness against jailbreak attacks has emerged as a significant concern. Zhao et al. (2024) pioneered the adversarial robustness evaluation of VLMs. Gong et al. (2023) introduce a benchmark called Safebench, which comprises 500 malicious typographies organized into 10 categories. Liu et al. (2024b) develop MM-SafetyBench, a benchmark featuring 5,040 text-image pairs across 13 scenarios. Additionally, Li et al. (2024b) compile the HADES-Dataset containing 750 harmful text-image pairs across 5 scenarios. Furthermore, Luo et al. (2024) propose a more comprehensive benchmark, JailBreakV-28K, which offers enhanced diversity and quality in harmful queries across 16 scenarios. Since this work focuses specifically on structure-based attack evaluation, we select Safebench, MM-SafetyBench and HADES-Dataset as the datasets for our experiments.
@@ 3 Threat Model
Since most of the state-of-the-art VLMs are only accessible via APIs, we follow the black-box attack framework (Gong et al., 2023). Under our setting, the attacker has no knowledge of the target model’s parameters or architecture and can perform attacks only in a single round of dialogue without prior context. The attacker is limited to querying the model and adjusting a few restricted hyper-parameters, such as the maximum token and temperature. Notably, we do not alter or introduce any system message in all experiments.
@@ 4.2.1 Encryption
To reduce the exposure of malicious content, we employ four strategies to encrypt images: word replacement, image mirroring, image rotation and base64 encoding.
@@
The base64 encoding method (Wei et al., 2023a; Handa et al., 2024), is to first encode the malicious query in base64 format then transform it into a typographical image.
@@ 4.3 Evil Alignment
Another limitation of existing methods is the lack of stealthy malicious guidance in text prompt. Since text safety alignment is more refined (Gong et al., 2023; Liu et al., 2024b), current structure-based attacks rely solely on neutral text prompts to guide the model in generating content related to images containing malicious information (Figure 1(a)). As a result, even when the model doesn’t directly refuse to respond, it often fails to generate the desired malicious content. Instead, it typically provides legal or ethical warnings and suggestions on how to avoid harmful behavior. Our experiments (Section 5.3) also confirm this.
@@
We refer to this as evil alignment and we empirically find it has a strong complementary effect to the encryption-decryption process. The complete prompts are included in Appendix A.3.
@@ 5.1 Setup
Table 1: Attack Success Rate (ASR) of baseline methods and MML (ours). FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024b). MML-$X$ represents different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. Best results are highlighted in bold. All evaluations are conducted without any system prompt.
@@
We conduct the experiments on three datasets: SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b) and HADES-Dataset (Li et al., 2024b), which are widely used as benchmarks for structure-based attacks. SafeBench includes 10 AI prohibited topics, selected based on the OpenAI Usage Policy (OpenAI, 2024d) and Meta’s Llama-2 Usage Policy (Meta, 2023). 50 malicious queries are generated by GPT-4 (OpenAI et al., 2024) for each topic, a total of 500 queries. MM-SafeBench comprises malicious queries across 13 scenarios. We filter out no-violation queries by using GPT-4o as moderation, getting a subset of 1,180 queries. HADES-Dataset (Li et al., 2024b) contains 750 malicious instructions across five scenarios. Further details can be found in Appendix A.4.
@@
We set FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b) as baseline methods for SafeBench and MM-SafeBench. For HADES-Dataset, we use HADES (Li et al., 2024b) as the baseline method. These methods all represent state-of-the-art structure-based or combination-based attacks. The Figstep and QueryRelated attacks are shown in Figure 1(a). HADES adds adversarial noise to the image similar to QueryRelated. For more details, please refer to the Appendix A.4.
@@
Previous methods have demonstrated high jailbreak success rates on VLMs such as LLaVA-1.5 (Liu et al., 2024a) and MiniGPT-4 (Chen et al., 2023), but they struggle with more advanced models like GPT-4o (OpenAI, 2024b). Therefore, we focuse on targeting these state-of-the-art models. The models included in our experiment are: GPT-4o-2024-08-06 (OpenAI, 2024b), GPT-4o-Mini-2024-07-18 (OpenAI, 2024a), Qwen-VL-Max-0809 (Bai et al., 2023), and Claude-3.5-Sonnet-20241022 (Anthropic, 2024), all of which represent the most advanced VLMs currently available (Contributors, 2023). The temperature of all models is set to 0.7, and we do not introduce any system messages for all experiments.
@@
Figure 4: ASR of baselines vs. MML-M (ours) across various topics in SafeBench. The left two figures presents the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b), while the right figure illustrates the ASR of MML using image mirroring as encryption method.
Figure 5: ASR of baselines vs. MML-M (ours) across various topics in $M$-SafeBench. The left two figures presents the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b), while the right figure illustrates the ASR of MML using image mirroring as encryption method.